# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v2

@@ 10.2 Frequency initialization
Like grid cells that fire at different frequencies [Hafting2005], our formalisms uses rotations at different scales, scale whose initialization has a major impact on position representations.
@@
Figure 6: (a) Distribution histogram of position in 1D navigation vs oscillation at the highest (green) and lowest (orange) frequencies. The longest distances reached by the model are comparable with the length of $\omega_{min}$ cycles.(b-c) Rotation blocks at high and low frequencies.
@@
The highest frequency $\omega_{max}$ defines the finest granularity $\Delta_{min}=\frac{2\pi}{\omega_{max}}$ at which the signal is encoded, while low frequencies encode long range dependencies. We would like the lowest frequency $\omega_{min}:=\frac{2\pi}{\Delta_{max}}$ to complete its cycle once the the model has covered the longest distance $\Delta_{max}$ it possibly can, where $\Delta_{max}=n$ on a $n$-size grid (fig. 6).
@@
Frequencies should encode signal at different scales, and a geometric initialization like the one used in RoPE [su2023roformerenhancedtransformerrotary] spreads the scales uniformly between 1 and $\Delta_{min}=\frac{1}{base}$ with:
@@
where $base\approx 10000$ in RoPE, dictates how many tokens the model has to see before completing a full circle. Compared to RoPE that always rotates by a constant angle $\Delta=1$, our input-dependent angles $\Delta_{t}$ can take values bellow $1$, and as such we might want our highest frequency $\omega_{max}$ to be greater than $1$. We extend the logic such that:
@@
10.3 3D and 5D navigation
@@
10.4 Groups beyond 2D rotations
So far, all our experiments were run with $2\times 2$ block-diagonal matrices, which we demonstrated in section 10.3 to be enough to capture translations in higher dimensions, as rotations on orthogonal circles. However, it is unlikely that the brain performs perfect rotations, and a model should be able to converge to the desired matrix representation. In this section, we decided to explore matrices beyond $2\times 2$ rotations, and analyze the tradeoff between model’s expressivity and inference speed. We will start by analyzing the impact of increasing the block size of our rotation matrices, as well as changing their structure, to not be skew-symmetric anymore. Finally, we will analyze non-commutative group structures, and show that they can be learned, at the cost of losing parallel sequence abilities.
@@ 10.4 Groups beyond 2D rotations
10.4.1 Other commutative groups
@@
Aside from block size, as Lie-group theory predicts, the nature of matrix $A$ influences the learned structure. The special linear group $\mathrm{SL}(n,\mathbb{R})$ – which preserves determinant, i.e. areas – can be represented using matrices with a null trace: $A=\overline{A}-\mathrm{Tr}(\overline{A})\implies\mathrm{Tr}(A)=0$. This group is more expressive than $SO(n)$ and in particular, we show that it can also be used to solve navigation tasks (fig. 7(a), MapEM-SLN, block-size 2, green).
Finally, we tested MapEM-NC, a non-commutative model that we detail further in sec. 10.4.2. This model learns a basis of rotations per block $S_{i},1\leq i\leq K$, allowing it to switch rotation axes on the fly. We show that such a model with block size $b=4$ and basis size $K=6$ (red) performs a par with MapEM-b2, and much better than a model with the same block size but a single generator matrix $S$ (MapEM-b4). This implies that what matters is the amount of rotations to perform rather than the block size in itself.
10.4.2 Non commutative Lie Groups
The ability to learn non commutative rules is fundamental, since rules as simple as "mother" and "father" should not commute. Indeed, the father of my mother is my grand father on my mother’s side, while the mother of my father is my grand mother on my father’s side, hence ending in a different position within the family tree/graph. This implies that the learned matrices do not commute, i.e.: $\mathbf{W}_{mother}\mathbf{W}_{father}\neq\mathbf{W}_{father}\mathbf{W}_{mother}$.
@@ 10.4.2 Non commutative Lie Groups
We introduce non-commutative models (NC), with a basis of K=n(n1)/2K=n(n-1)/2 skew-symmetric matrices $\{S_{i}\}_{1\leq i\leq K}$, $\theta:=\Delta_{t}\omega\in\mathbb{R}^{n_{h}\times n_{b}\times K}$ and $R_{\theta}=\exp\left(\sum_{i=1}^{K}\theta_{i}S_{i}\right)$. The non commutativity forces us to compute path-integration sequentially and use a parallel scan to speed-up computation, since our models cannot leverage the parallel processing abilities of Transformers anymore, making them analogous to TEM-t [whittington2022relatingtransformersmodelsneural], where path integration is computed sequentially before computing attention.
@@
This section covers a non exhaustive list of possibilities that our formalism offers, even though future research should help to understand the ideal tradeoff between inference speed and models expressivity.
10.5 Attention maps
To give a visual intuition of the increased recall of MapEM-s models, we plot in fig 8 the attention maps of four models when trying to retrieve a memory at a previously visited locations (highlighted in red). As expected, RoPE’s attention is noisy and the model never manages to find the correct memory. Models who’s attention rely on both structure and observation, like MapWM and MapEM-os manage to find the appropriate memory, but their attention map is not as sparse as the one where position alone is used, showing that structural attention $A_{P}$ acts as an attention mask to ignore structurally irrelevant tokens. Since positions vary smoothly from one token to the next, so should their features $p_{t}$, explaining why the attention values for the MapEM-s model are rarely above $0.5$, but can still differentiate tokens efficiently. Note that note all positions on the grid have an observation, explaining why certain positions have a high similarity (looped made by the model) but no red dot on them since there’s no recall to be made.
(a) RoPE
(b) MapWM
(c) MapEM-os
(d) MapEM-s
Figure 8: MapFormers learn sparse attention maps. RoPE / WM / EM attention maps. Red dots designate objects at previously visited location. (a) RoPE fails to attend to the correct token. (b-c) MapWM and MapEM-os models manage to attend to the relevant token but (d) MapEM-s exhibits a sparse attention map as it only needs to focus on abstract position $p_{t}$, while keys and queries can repeat an increase the similarity of structurally distinct tokens

### v3

@@ A.8 Angular velocities initialization
Like grid cells that fire at different frequencies (Hafting et al., 2005), our formalisms uses rotations at different scales, scale whose initialization has a major impact on position representations.
@@
Figure 6: (a) Distribution histogram of position in 1D navigation vs oscillation at the highest (green) and lowest (orange) angular velocities. The longest distances reached by the model are comparable with the length of $\omega_{min}$ cycles.(b-c) Rotation blocks at high and low frequencies.
@@
The highest angular velocity (or frequency $f$, assuming $\omega=2\pi f$) $\omega_{max}$ defines the finest granularity $\Delta_{min}=\frac{2\pi}{\omega_{max}}$ at which the signal is encoded, while low frequencies encode long range dependencies. We would like the lowest angular velocity $\omega_{min}:=\frac{2\pi}{\Delta_{max}}$ to complete its cycle once the the model has covered the longest distance $\Delta_{max}$ it possibly can, where $\Delta_{max}=n$ on a $n$-size grid (fig. 6).
@@
Angular velocities should encode signal at different scales, and a geometric initialization like the one used in RoPE (Su et al., 2023) spreads the scales uniformly between 1 and $\Delta_{min}=\frac{1}{base}$ with:
@@
where $base\approx 10000$ in RoPE, dictates how many tokens the model has to see before completing a full circle. Compared to RoPE that always rotates by a constant rotation duration $\Delta=1$, our input-dependent rotation durations $\Delta_{t}$ can take values bellow $1$, and as such we might want our highest angular velocity $\omega_{max}$ to be greater than $1$. We extend the logic such that:
@@
Appendix B Further experiments and methods
When generating trajectories, to favor grid exploration, we uniformly sample a direction and the number of steps $k$ to pursue in that direction $(1\leq k\leq 10)$. Otherwise, agents would be statistically stuck in small environments.
For 1D/2D grid navigation, we trained all models on sequences of $l=128$ steps, a grid size of $l_{grid}=64$, and a probability of placing an object at a specific location $p_{empty}=0.5$. We tested length generalization on two OOD datasets - OOD-d: $l=64,p_{empty}=0.2,l_{grid}=32$ and OOD-s: $l=512,p_{empty}=0.8,l_{grid}=128$.
Unless stated otherwise, for 1D-2D navigation and beyond, we trained single layer models with two heads of size $h=64$ on 200K sequences, with AdamW optimizer (Loshchilov and Hutter, 2019), a linear learning rate decay, a base learning rate $lr=3e^{-4}$, weight decay $0.05$ and a batch size of $128$ on a single GPU.
B.1 3D and 5D navigation
@@ Appendix B Further experiments and methods
B.2 Non-commutative groups
So far, all our experiments were run with $2\times 2$ block-diagonal matrices, which we demonstrated in section B.1 to be enough to capture translations in higher dimensions, as rotations on orthogonal circles. However, it is unlikely that the brain performs perfect rotations, and a model should be able to converge to the desired matrix representation. In this section, we decided to explore matrices beyond $2\times 2$ rotations, and analyze the tradeoff between model’s expressivity and inference speed. We will start by analyzing the impact of increasing the block size of our rotation matrices, as well as changing their structure, to not be skew-symmetric anymore. Finally, we will analyze non-commutative group structures, and show that they can be learned, at the cost of losing parallel sequence abilities.
@@ B.2 Non-commutative groups
B.2.1 Other commutative groups
@@
Aside from block size, as Lie-group theory predicts, the nature of matrix $A$ influences the learned structure. The special linear group $\mathrm{SL}(n,\mathbb{R})$ – which preserves determinant, i.e. areas – can be represented using matrices with a null trace: $A=\overline{A}-\mathrm{Tr}(\overline{A})\implies\mathrm{Tr}(A)=0$. This group is more expressive than $SO(n)$ (but not compact), and we show that it can also be used to solve navigation tasks (fig. 7(a), MapEM-SLN, block-size 2, green).
Finally, we tested MapEM-NC, a non-commutative model that we detail further in sec. B.2.2. This model learns a basis of rotations per block $S_{i},1\leq i\leq K$, allowing it to switch rotation axes on the fly. We show that such a model with block size $b=4$ and basis size $K=6$ (red) performs a par with MapEM-b2, and much better than a model with the same block size but a single generator matrix $S$ (MapEM-b4). This implies that what matters is the amount of rotations to perform rather than the block size in itself.
B.2.2 Non commutative Lie Groups
The ability to learn non commutative rules is fundamental, since rules as simple as ”mother” and ”father” should not commute. Indeed, the father of my mother is my grand father on my mother’s side, while the mother of my father is my grand mother on my father’s side, hence ending in a different position within the family tree/graph. This implies that the learned matrices do not commute, i.e.: $\mathbf{W}_{mother}\mathbf{W}_{father}\neq\mathbf{W}_{father}\mathbf{W}_{mother}$.
@@ B.2.2 Non commutative Lie Groups
We introduce non-commutative models (NC), with a basis of K=n(n1)/2K=n(n-1)/2 skew-symmetric matrices $\{S_{i}\}_{1\leq i\leq K}$, $\theta:=\Delta_{t}\omega\in\mathbb{R}^{n_{h}\times n_{b}\times K}$ and $R_{\theta}=\exp\left(\sum_{i=1}^{K}\theta_{i}S_{i}\right)$. The non commutativity forces us to compute path-integration sequentially and use a parallel scan to speed-up computation, since our models cannot leverage the parallel processing abilities of Transformers anymore, making them analogous to TEM-t (Whittington et al., 2022), where path integration is computed sequentially before computing attention.